OpenStudy (anonymous):

Simplify the permutation: P(10, 4) 5040 720 24 240

OpenStudy (anonymous):

I think it should be 5040. What it is saying is that there are 10 possible choices, and you want only four of those ten, so what you would do is 10!/(10-4)!, which gives you 5040.

OpenStudy (apoorvk):

\[P(n,r) = ~^nP_r\] \[^nP_r = \frac{n!}{(n-r)!}\] Use this^ @ladynisha

OpenStudy (anonymous):

nPr n = 10 r = 4 nPr = n!/(n - r)! 10P4 = 10!/(10 - 4)! 10P4 = 10!/6! 10P4 = 10*9*8*7*6!/6! 10P4 = 10*9*8*7 = 5040 :)

OpenStudy (anonymous):

THANKS